# SwiftSeg: Efficient Training-Free Open-Vocabulary Segmentation via Hierarchical Attention Refinement Method

### v2

@@ Abstract
Open-vocabulary semantic segmentation (OVSS) aims to segment objects from arbitrary text categories without requiring densely annotated datasets. Although contrastive learning based models enable zero-shot segmentation, they often lose fine spatial precision at pixel level, due to global representation bias. In contrast, diffusion-based models naturally encode fine-grained spatial features via attention mechanisms that capture both global context and local details. However, they often face challenges in balancing the number of iterations with the quality of the segmentation. In this work, we propose SwiftSeg, a novel and efficient training-free framework with only (1+1)-step of reverse process of a pretrained diffusion model. Moreover, instead of running multiple times for different classes, SwiftSeg performs segmentation for all classes at once. To further enhance the segmentation quality, SwiftSeg introduces three key components: (i) a dual-prompt mechanism for discriminative, class-aware attention extraction, (ii) a Hierarchical Attention Refinement Method (HARD) that enhances fused cross-attention using scale-aligned self-attention maps, and (iii) a Test-Time Flipping (TTF) scheme designed to improve spatial consistency. Extensive experiments show that SwiftSeg achieves state-of-the-art training-free performance, obtaining 43.8% average mIoU across PASCAL VOC, PASCAL Context, and COCO Object benchmarks while maintaining superior inference efficiency. Our results demonstrate that SwiftSeg provides a strong foundation for extendability, bridging the gap between segmentation quality and inference efficiency.
@@
Open-vocabulary segmentation , Diffusion models , Training-free , Attention Refinement
@@ 1 Introduction
Figure 1: Open-vocabulary semantic segmentation results using SwiftSeg. We visualize predictions on challenging scenarios, including contextual images, camouflaged objects, and internet-collected images. The highlighted results demonstrate that, in addition to accurate segmentation of both stuff and things, our method effectively handles camouflaged objects, distinguishes between different characters or dog breeds, and differentiates objects based on color variations.
Existing OVSS methods primarily follow two dominant paradigms: vision-language models and diffusion models. Vision-language models such as CLIP[1] leverage contrastive learning to align visual and textual representations within a shared embedding space. Approaches based on CLIP can be broadly categorized into training-free methods [2, 3, 4, 5], which directly utilize pre-trained models, and training methods [6, 7, 8, 9, 10, 11, 12, 13, 14], which improve multimodal alignment through additional supervision. However, these models often suffer from limited spatial precision and struggle to accurately capture object boundaries due to their reliance on global image-text embeddings. In contrast, diffusion models [15, 16, 17, 18, 19, 20, 21] have recently emerged as a powerful alternative, as their generative attention mechanisms inherently preserve spatial structures, enabling them to capture both object location and shape with high fidelity. Nevertheless, many diffusion-based methods require extensive supervised training and computationally expensive synthetic dataset generation, limiting their scalability and practical deployment.
@@
In this paper, we propose SwiftSeg, an efficient zero-shot framework for open-vocabulary semantic segmentation (OVSS) that leverages text-to-image diffusion models, as shown in Figure 1. SwiftSeg balances high segmentation quality with computational efficiency through key innovations. It performs image reconstruction using DDIM Inversion, enabling the simultaneous extraction of both self-attention and cross-attention maps across all candidate classes. A dual-prompt mechanism—a text prompt for reconstruction and a class prompt for generating class-specific cross-attention maps—enables multi-class segmentation without repeated inference. Additionally, SwiftSeg introduces the Hierarchical Attention Refinement Method (HARD), which aggregates attention maps from multiple scales, combining fine details from higher resolutions and broader context from lower resolutions to improve localization, shape accuracy, and boundary delineation. Finally, the Test-Time Flipping (TTF) strategy enhances segmentation accuracy by averaging predictions from both the original and flipped images, improving spatial consistency and robustness.
We summarize our key contributions as follows:
  • 1.
    We introduce SwiftSeg, a novel training-free open-vocabulary segmentation method that leverages diffusion models with fast (1+1)-step DDIM inversion and a unified class prompt, enabling efficient attention map generation for all candidate classes in a single inference run.
  • 2.
    To enhance segmentation accuracy, we propose three key techniques: (i) a dual-prompt mechanism that separately guides semantic reconstruction and class-specific attention extraction, (ii) HARD proposed that improves fused cross-attention maps using transformed self-attention maps, and (iii) TTF that improves the spatial consistency of the segmentation mask.
  • 3.
    We summarize our key contributions as follows: We conduct comprehensive experiments on three benchmark datasets—PASCAL VOC, PASCAL Context, and COCO Object—demonstrating that SwiftSeg achieves state-of-the-art performance among training-free methods while maintaining superior computational efficiency.
@@ 2.1 Contrastive Vision-Language Models for OVSS
In this study, we propose a training-free open-vocabulary semantic segmentation (OVSS) approach, driven by the goal of reducing annotation overhead and enhancing scalability—while maintaining flexibility in segmenting diverse. Unlike prior methods that predominantly rely on vision-language models with contrastive alignment, which often compromise spatial precision and object boundary accuracy, our approach adopts a generative pre-trained model. By inherently preserving the spatial structure of diffusion models through attention maps, we enable more accurate and semantically grounded segmentation without the need for additional training or pixel-level annotations.
@@ 2.2 Semantic Segmentation with Diffusion Models
Figure 2: Overview of the proposed SwiftSeg. he input consists of both the original and horizontally flipped images, which are first reconstructed via DDIM inversion to obtain their latent representations. In parallel, a text prompt $\mathcal{P}$ and a class prompt $\mathcal{P^{\prime}}$ are generated to guide inversion and extract class-specific cross-attention maps corresponding to candidate classes (e.g., bus, motorbike, sheep). Cross-attention and self-attention maps at multiple resolutions are then extracted, transformed, and fused via weighted aggregation, where the self-attention maps are used to refine the class-specific cross-attention maps. Final segmentation masks are derived from the refined per-class score maps. Test-Time Flipping (TTF) is applied during inference to enhance robustness: attention maps from the flipped image are spatially realigned and averaged with those from the original image to improve prediction reliability.
@@ 3 Methods
3.1 SwiftSeg Overview
Our objective is to perform training-free open-vocabulary semantic segmentation, which assigns each pixel to one of the candidate categories $\mathcal{C}=\{c_{1},c_{2},...,c_{M}\}$ without requiring dense pixel-level annotations. Given a real image $I_{0}$ and a candidate class set $\mathcal{C}$, our proposed SwiftSeg framework extracts segmentation masks via cross-attention and self-attention analysis from a pre-trained Stable Diffusion (SD) model. The overall framework is depicted in Figure 2. The entire pipeline consists of three stages:
  • 1.
    Stage 1: Prompt Construction (Sec. 3.2): We first construct two complementary text prompts: the caption prompt $\mathcal{P}$, and the class prompt $\mathcal{P}^{\prime}$ containing the candidate categories. This dual-prompt enables the diffusion model to simultaneously reconstruct the image while injecting class-awareness into attention maps.
  • 2.
    Stage 2: Attention Extraction via Efficient Inversion (Sec. 3.3): We employ a fast (1+1)-step DDIM inversion using 2-Rectified Flow to reconstruct the input image while extracting both cross-attention $\mathcal{C}^{r}$ and self-attention $\mathcal{S}^{r}$ maps across multiple resolutions $r\in\mathcal{R}$. Cross-attention captures class-to-region alignment, while self-attention models intra-image structure.
  • 3.
    Stage 3: Hierarchical Attention Refinement (Sec. 3.4): To produce accurate segmentation, cross-attention maps across multiple resolutions are fused into $\widehat{\mathcal{C}}_{c}$ via weighted aggregation. The fused cross-attention is then further refined using affinity matrices derived from self-attention across resolutions. Finally, we apply Test-Time Flipping (TTF) to enhance spatial consistency and produce the final segmentation masks.
3.2 Prompt Construction
@@
To enable semantic reconstruction and class-aware attention extraction, SwiftSeg introduces a dual-prompt mechanism composed of a textual prompt $\mathcal{P}$ and a class prompt $\mathcal{P}^{\prime}$. The textual prompt $\mathcal{P}$, generated by an image captioning model (e.g., BLIP), serves as the text condition for image reconstruction via DDIM inversion, facilitating faithful image generation. To explicitly inform the model about candidate classes, we append the class prompt $\mathcal{P}$ to the textual prompt, following the structure: "image caption; $\mathcal{P}$’". This design enables the diffusion model to focus on class-specific semantics during generation. The class prompt $\mathcal{P}^{\prime}$ is constructed from a candidate label set $\mathcal{C}=\{c_{1},\ldots,c_{M}\}$, containing of $M$ open-vocabulary categories. For example, if $\mathcal{C}$ = { tv monitor, dog, person }, then $\mathcal{P}^{\prime}=``tv{\color[rgb]{1,1,1}s}monitor,$ $dog,$ $person"$. These candidate labels are essential for computing class-specific cross-attention maps that connect textual concepts to spatial regions in the semantic reconstruction of the image. Unlike contrastive models such as CLIP, which inherently suppress irrelevant concepts while enhancing relevant ones, text-to-image models like Stable Diffusion (SD) do not automatically ignore out-of-context objects. Instead, SD can preserve unrelated visual concepts as semantic noise. This behavior is illustrated in Figure 3, where distractor classes (those not present in the image) still produce activated attention maps. This motivates the need for a tailored class candidate set $\mathcal{C}$ for each image.
@@ 3.3.1 Preliminaries
In text-conditioned diffusion models, such as Stable Diffusion [39], generate an image by denoising a latent variable $z_{T}$ under the condition of a text prompt $P$. To enable this process, the network $\epsilon_{\theta}$ is trained to predict the noise added during the forward diffusion process. The objective is to minimize the following loss function:
@@
$z_{t-1}$ $=\sqrt{\alpha_{t-1}}\left[\frac{z_{t}-\sqrt{1-\alpha_{t}}\,\epsilon_{\theta}\bigl{(}z_{t},t,\tau_{\theta}(\mathcal{P})\bigr{)}}{\sqrt{\alpha_{t}}}\right]$ (2)
@@ 3.3.2 Latent Inversion for Semantic Reconstruction
Since DDIM inversion involves both forward and reverse passes, its computational complexity scales with $O(2T)$, posing a trade-off between reconstruction fidelity and inference time. To overcome this bottleneck, we employ 2-Rectified Flow (2-RF) [43], a distilled diffusion model trained to perform a faithful inversion with minimal steps. 2-RF straightens the latent space ODE trajectories via a reflow procedure, enabling accurate inversion in a single run. Utilizing 2-RF allows accurate inversion with only two steps $O(2)$, yielding significant speed-up while preserving reconstruction quality. By using Test-Time Flipping (TTF), we perform the reconstruction process for both the original image and the flipped image, with a batch size of 2.
@@ 3.3.3 Attention Extraction for Segmentation
SwiftSeg generates semantic segmentation masks by extracting cross-attention and self-attention maps during the Stable Diffusion denoising process (latent $z_{1}$ to $z_{0}$). The cross-attention maps $\mathcal{C}^{l,r}\in[0,1]^{r\times r\times d_{\mathcal{P}}}$ capture spatial correlations between text tokens and image regions. On the other hand, self-attention maps $\mathcal{S}^{l,r}\in[0,1]^{r\times r\times r\times r}$ preserve the spatial structure of the image by modeling pairwise similarities between image positions. These attention maps are computed across multiple resolutions =$\mathcal{R}=\{8,16,32,64\}$, with each resolution having $L$ layers. The maps for each resolution are averaged over layers to obtain a final representation:
$\mathcal{C}^{r}=\frac{1}{L}\sum_{l=1}^{L}\text{Softmax}\left(\frac{Q_{z}K_{\tau_{\theta}(\mathcal{P})}^{\top}}{\sqrt{d}}\right).$ (6)
$\mathcal{S}^{r}=\frac{1}{L}\sum_{l=1}^{L}\text{Softmax}\left(\frac{Q_{z}K_{z}^{\top}}{\sqrt{d}}\right)$ (7)
where $Q_{z}$, $K_{z}$, and $K_{\tau_{\theta}(\mathcal{P})}$ represent the query and key matrices derived from linear projections. Unlike most prior methods that generate segmentation masks for each candidate class independently, requiring $M$ inference steps for $M$ candidate classes, we propose a method that generates segmentation masks for all candidate classes simultaneously in a single inference run. By using a class prompt $\mathcal{P}^{\prime}$ we control the cross-attention for the candidate classes rather than relying solely on the text prompt $\mathcal{P}$. While $\mathcal{P}$ continues to guide image reconstruction via DDIM inversion, $\mathcal{P}^{\prime}$ enables focused cross-attention maps for the candidate classes and offers greater flexibility in specifying candidate classes compared to the text prompt. In this setting, cross-attention maps $\mathcal{C}^{l}$ in Equation 6 are computed using $K_{\tau_{\theta}(\mathcal{P^{\prime}})}$ instead of $K_{\tau_{\theta}(\mathcal{P})}$ This approach enables the generation of cross-attention maps for all candidate classes within a single inference run.
@@ 3.4 Hierarchical Attention Refinement Method
In semantic segmentation tasks, attention maps play a crucial role in localizing and identifying different regions within an image. However, attention maps at a fixed resolution provide valuable localization information but may fail to capture the semantic diversity present across multiple spatial scales. As illustrated in Figure 5(a), cross-attention maps at various resolutions offer complementary insights: $\mathcal{C}^{8\times 8}$ provides coarse maps but accurate localization, while $\mathcal{C}^{16\times 16}$ captures better details of object position and shape. At higher resolutions (e.g., $\mathcal{C}^{32\times 32}$ and $\mathcal{C}^{64\times 64}$) attention maps may become increasingly blurred and noisy. These observations align with previous studies [44], which show that higher-resolution layers often focus on superficial features (such as color or style), while lower-resolution layers encode object semantics. To address this limitation, we using a hierarchical cross attention fusion mechanism that emphasizes semantically cross attention maps (e.g., $\mathcal{C}^{16\times 16}$) while down-weighting noisy maps (e.g., $\mathcal{C}^{32\times 32}$ or $\mathcal{C}^{64\times 64}$). The fused cross-attention map $\widehat{\mathcal{C}}_{c}$ is computed as a weighted sum of the resolution-specific maps, each upsampled to a size of $64\times 64$:
@@
Each self-attention tensor $\mathcal{S}^{r}\in[0,1]^{r\times r\times r\times r}$ captures pairwise affinities between all locations in the latent representation at resolution $r$. Specifically, for a given position $(i,j)$, we extract a 2D attention map $\mathcal{S}^{r}_{[i,j,:,:]}\in[0,1]^{r\times r}$, which reflects the semantic context of the object group in the corresponding region of the original image. Figure 6 visualizes self-attention maps at different resolutions and spatial locations. The resolution of each attention map determines its receptive field size relative to the original image.
To leverage self-attention information into the segmentation process, we refine the fused cross-attention map $\widehat{\mathcal{C}_{c}}$ by applying affinity weighting derived from the self-attention maps. Since $\widehat{\mathcal{C}_{c}}\in R^{64\times 64}$ is defined at a fixed spatial resolution, we must transform the self-attention tensors $\mathcal{S}^{r}\in R^{r\times r\times r\times r}$ to $\mathcal{S}^{r\rightarrow 64}\in R^{64\times 64\times 64\times 64}$ through upsampling and repetition. An example of this transformation from. An example of transformation from size ($2\times 2\times 2\times 2$) to ($4\times 4\times 4\times 4$) is shown in Figure 7. While upsampling the last two dimensions of the 4D tensor (which consists of 2D tensors) is straightforward, the repetition operation on the first two dimensions is based on empirical observations: 2D attention maps at neighboring positions often correspond to the same object group. For example, $\mathcal{S}^{r}_{[i,j,:,:]}$ is similar to $\mathcal{S}^{r}_{[i+1,j,:,:]}$ or $\mathcal{S}^{r}_{[i,j+1,:,:]}$, particularly when they represent the same object region.
This transformation produces spatially aligned 4D self-attention tensors $\mathcal{S}^{r\rightarrow 64}\in[0,1]^{64\times 64\times 64\times 64}$, which are compatible for multiplicative refinement of the class-specific cross-attention map $\widehat{\mathcal{C}}_{c}\in[0,1]^{64\times 64}$. To incorporate self-attention affinities at multiple resolutions $r\in\mathcal{R}$, each $\mathcal{S}^{r\rightarrow 64}$ is matricized into an affinity matrix, and $\widehat{\mathcal{C}}_{c}$ is vectorized to match this dimension. These maps are then normalized and combined via a weighted sum over all resolutions to yield the final refined segmentation mask:
@@
By calculating $\widehat{\mathcal{A}}$ for the entire candidate class list $C$, we obtain the set of segmentation maps $A=\{\widehat{\mathcal{A}}_{c}\mid c\in C\}$. Then, by applying the Test-Time Flipping (TTF) technique, we obtain two corresponding segmentation map lists: $\mathbf{A}^{\text{origin}}$ from the original image and $\mathbf{A}^{\text{flipped}}$ from the horizontally flipped image. We align the spatial information of the attention maps from the flipped image to match those from the original image. Then, we compute the average of each pair of corresponding attention maps from both lists to improve spatial consistency and prediction accuracy.
$\widehat{\mathcal{A}}_{c}^{*}=\frac{1}{2}\left[\widehat{\mathcal{A}}_{c}^{\text{origin}}+FLIP\Bigl{(}\widehat{\mathcal{A}}_{c}^{\text{flipped}}\Bigl{)}\right]$

### v3

# Writing – review and editing, Supervision, Conceptualization

@@ Abstract
Open-vocabulary semantic segmentation (OVSS) aims to segment objects from arbitrary text categories without requiring densely annotated datasets. Although contrastive learning based models enable zero-shot segmentation, they often lose fine spatial precision at pixel level, due to global representation bias. In contrast, diffusion-based models naturally encode fine-grained spatial features via attention mechanisms that capture both global context and local details. However, they often face challenges in balancing the computation costs and the quality of the segmentation mask. In this work, we present FA-Seg, a Fast and Accurate training-free framework for open-vocabulary segmentation based on diffusion models. FA-Seg performs segmentation using only a (1+1)-step from a pretrained diffusion model. Moreover, instead of running multiple times for different classes, FA-Seg performs segmentation for all classes at once. To further enhance the segmentation quality, FA-Seg introduces three key components: (i) a dual-prompt mechanism for discriminative, class-aware attention extraction, (ii) a Hierarchical Attention Refinement Method (HARD) that enhances semantic precision via multi-resolution attention fusion, and (iii) a Test-Time Flipping (TTF) scheme designed to improve spatial consistency. Extensive experiments show that FA-Seg achieves state-of-the-art training-free performance, obtaining 43.8% average mIoU across PASCAL VOC, PASCAL Context, and COCO Object benchmarks while maintaining superior inference efficiency. Our results demonstrate that FA-Seg provides a strong foundation for extendability, bridging the gap between segmentation quality and inference efficiency. The source code will be open-sourced after this paper is accepted.
@@
[…]
@@ 1 Introduction
Figure 1: Open-vocabulary semantic segmentation results using FA-Seg. We visualize predictions on challenging scenarios, including contextual images, camouflaged objects, and internet-collected images. The highlighted results demonstrate that, in addition to accurate segmentation of both stuff and things, our method effectively handles camouflaged objects, distinguishes between different characters or dog breeds, and differentiates objects based on color variations.
Existing OVSS methods primarily follow two dominant paradigms: vision-language models and generative models. Vision-language models such as CLIP [1] leverage contrastive learning to align visual and textual representations within a shared embedding space. Approaches based on CLIP can be broadly categorized into training-free methods [2, 3, 4, 5], and training methods [6, 7, 8, 9, 10, 11, 12, 13, 14]. However, these models often suffer from limited spatial precision due to their reliance on global image-text embeddings. In contrast, generative models [15, 16, 17, 18, 19, 20, 21] have recently emerged as a powerful alternative, as their generative attention mechanisms inherently preserve spatial structures, enabling them to capture both object location and shape with high fidelity. Nevertheless, many diffusion-based methods require extensive supervised training and computationally expensive synthetic dataset generation, limiting their scalability and practical deployment.
@@
In this paper, we present FA-Seg, an efficient training-free framework for open-vocabulary semantic segmentation (OVSS) that leverages pretrained text-to-image diffusion models, as illustrated in Figure 1. FA-Seg reconstructs the input image using fast (1+1)-step DDIM inversion. The dual-prompt mechanism is used in addition to class-aware semantic enhancement, allowing for simultaneous multi-class segmentation without iterative inference. In addition to Test-Time Flipping (TTF), to further improve segmentation quality, we propose a multi-scale attention refinement strategy, HARD (Hierarchical Attention Refinement Design), which fuses cross-attention and self-attention maps across resolutions to enhance object localization and boundary accuracy. We summarize our key contributions as follows:
  • We introduce FA-Seg that leverages diffusion models with fast (1+1)-step DDIM inversion and a unified class prompt, enabling efficient attention map generation for all candidate classes in a single inference run.
  • To enhance segmentation accuracy, we propose three key techniques: (i) a dual-prompt mechanism that separately guides semantic reconstruction and class-specific attention extraction to help better class awareness, (ii) HARD proposed that improves fused cross-attention maps using transformed self-attention maps at multi-scales, and (iii) TTF that improves the spatial consistency of the segmentation mask.
  • We conduct comprehensive experiments on three benchmark datasets—PASCAL VOC, PASCAL Context, and COCO Object—demonstrating that FA-Seg achieves state-of-the-art performance among training-free methods while maintaining superior computational efficiency.
@@ 2.1 Contrastive Vision-Language Models for OVSS
In this study, we propose a training-free open-vocabulary semantic segmentation (OVSS) approach, driven by the goal of reducing annotation overhead and enhancing scalability—while maintaining flexibility in segmenting diverse. Unlike prior methods that predominantly rely on vision-language models with contrastive alignment, which often compromise spatial precision, our approach adopts a generative pre-trained model. By inherently preserving the spatial structure of diffusion models through attention maps, we enable more accurate and semantically grounded segmentation without the need for additional training or pixel-level annotations.
@@ 2.2 Semantic Segmentation with Diffusion Models
Figure 2: Overview of the proposed FA-Seg. he input consists of both the original and horizontally flipped images, which are first reconstructed via DDIM inversion to obtain their latent representations. In parallel, a text prompt $\mathcal{P}$ and a class prompt $\mathcal{P^{\prime}}$ are generated to guide inversion and extract class-specific cross-attention maps corresponding to candidate classes (e.g., bus, motorbike, sheep). Cross-attention and self-attention maps at multiple resolutions are then extracted, transformed, and fused via weighted aggregation, where the self-attention maps are used to refine the class-specific cross-attention maps. Final segmentation masks are derived from the refined per-class score maps. Test-Time Flipping (TTF) is applied during inference to enhance robustness: attention maps from the flipped image are spatially realigned and averaged with those from the original image to improve prediction reliability.
@@ 3 Methods
3.1 FA-Seg Overview
Our objective is to perform training-free open-vocabulary semantic segmentation without requiring dense pixel-level annotations. Given a real image $I_{0}$ and a candidate class set, our proposed FA-Seg framework extracts segmentation masks via cross-attention and self-attention analysis from a pre-trained Stable Diffusion (SD) model. The overall framework is depicted in Figure 2. The entire pipeline consists of three stages:
  • Stage 1: Dual-Prompt Construction (Sec. 3.2): We first construct two complementary text prompts: the caption prompt $\mathcal{P}$, and the class prompt $\mathcal{P}^{\prime}$ containing the candidate categories.
  • Stage 2: Attention Extraction via Efficient Inversion (Sec. 3.3): We employ a fast DDIM inversion using diffusion model to reconstruct the input image while extracting both cross-attention and self-attention maps across multiple resolutions $r\in\mathcal{R}$.
  • Stage 3: Hierarchical Attention Refinement (Sec. 3.4): To produce accurate segmentation, cross-attention maps across multiple resolutions are fused into $\widehat{\mathcal{C}}_{c}$ via weighted aggregation. The fused cross-attention is then further refined using affinity matrices derived from self-attention across resolutions.
3.2 Dual-Prompt Construction
@@
To enable semantic reconstruction and class-aware attention extraction, FA-Seg introduces a dual-prompt mechanism composed of $\Bigl{(}\mathcal{P}$, $\mathcal{P}^{\prime}\Bigl{)}$. The textual prompt $\mathcal{P}$, generated by an image captioning model (e.g., BLIP), serves as the text condition for image reconstruction via DDIM inversion, facilitating faithful image generation. To explicitly inform the model about candidate classes, we append the class prompt $\mathcal{P}$ to the textual prompt, following the structure: "image caption; $\mathcal{P}$’". This design enables the diffusion model to focus on class-specific semantics during generation. The class prompt $\mathcal{P}^{\prime}$ is constructed from a candidate set $\mathcal{C}=\{c_{1},\ldots,c_{M}\}$. For example, if $\mathcal{C}$ = { tv monitor, dog, person }, then $\mathcal{P}^{\prime}=``tv{\color[rgb]{1,1,1}s}monitor,$ $dog,$ $person"$. These candidate labels are essential for computing class-specific cross-attention maps that connect textual concepts to spatial regions in the semantic reconstruction of the image. Unlike contrastive models such as CLIP, which inherently suppress irrelevant concepts while enhancing relevant ones, text-to-image models like Stable Diffusion (SD) do not automatically ignore out-of-context objects. Instead, SD can preserve unrelated visual concepts as semantic noise. This behavior is illustrated in Figure 3, where distractor classes (those not present in the image) still produce activated attention maps. This motivates the need for a tailored class candidate set $\mathcal{C}$ for each image.
@@
To construct the candidate set $\mathcal{C}$, we adopt the approach from previous works [20, 21], which use BLIP and CLIP to identify potential object categories from a predefined vocabulary in images. To enhance candidate set quality, we use CLIP with TagCLIP — a variant of CLIP tailored for multi-label classification. This improved pairing enables more accurate category predictions, resulting in a more reliable candidate set $\mathcal{C}$.
@@ 3.3.1 Preliminaries
In text-conditioned diffusion models, such as Stable Diffusion [39], generate an image by denoising a latent variable $z_{T}$ under the condition of a text prompt $\mathcal{P}$. To enable this process, the network $\epsilon_{\theta}$ is trained to predict the noise added during the forward diffusion process. The objective is to minimize the following loss function:
@@
$z_{t-1}$ $=\sqrt{\alpha_{t-1}}\left[\frac{z_{t}-\sqrt{1-\alpha_{t}}\,\epsilon_{\theta}\bigl{(}z_{t},t,\tau_{\theta}(\mathcal{P})\bigr{)}}{\sqrt{\alpha_{t}}}\right]$
@@ 3.3.2 Latent Inversion for Semantic Reconstruction
Since DDIM inversion involves both forward and reverse passes, its computational complexity scales with 2T steps, posing a trade-off between reconstruction fidelity and inference time. To overcome this bottleneck, we employ 2-Rectified Flow (2-RF) [43], a distilled diffusion model trained to perform a faithful inversion with minimal steps. 2-RF straightens the latent space ODE trajectories via a reflow procedure, enabling accurate inversion in a single run. Utilizing 2-RF allows accurate inversion with only two steps or (1+1)-step, yielding significant speed-up while preserving reconstruction quality.
@@ 3.3.3 Attention Extraction for Segmentation
FA-Seg generates semantic segmentation masks by extracting cross-attention and self-attention maps during the Stable Diffusion denoising process (latent $z_{1}$ to $z_{0}$). The cross-attention maps $\mathcal{C}^{l,r}\in[0,1]^{r\times r\times d_{\mathcal{P}}}$ capture spatial correlations between text tokens and image regions. On the other hand, self-attention maps $\mathcal{S}^{l,r}\in[0,1]^{r\times r\times r\times r}$ preserve the spatial structure of the image by modeling pairwise similarities between image positions. These attention maps are computed across multiple resolutions =$\mathcal{R}=\{8,16,32,64\}$, with each resolution having $L$ layers. The maps for each resolution are averaged over layers to obtain a final representation:
$\mathcal{C}^{r}=\frac{1}{L}\sum_{l=1}^{L}\text{Softmax}\left(\frac{Q^{l}_{z}K_{\tau_{\theta}(\mathcal{P})}^{\top}}{\sqrt{d}}\right).$ (5)
$\mathcal{S}^{r}=\frac{1}{L}\sum_{l=1}^{L}\text{Softmax}\left(\frac{Q^{l}_{z}K_{z}^{\top}}{\sqrt{d}}\right)$ (6)
where $Q_{z}$, $K_{z}$, and $K_{\tau_{\theta}(\mathcal{P})}$ represent the query and key matrices derived from linear projections. By using a class prompt $\mathcal{P}^{\prime}$ we control the cross-attention for the candidate classes rather than relying solely on the text prompt $\mathcal{P}$. While $\mathcal{P}$ continues to guide image reconstruction via DDIM inversion, we propose using $\mathcal{P}^{\prime}$ to generates segmentation masks for all candidate classes and offers greater flexibility in specifying candidate classes compared to the text prompt $\mathcal{P}$. This dual-prompt enables the diffusion model to simultaneously reconstruct the image while injecting class-awareness into attention maps. In this setting, cross-attention maps $\mathcal{C}^{l}$ in Equation 5 are computed using $K_{\tau_{\theta}(\mathcal{P^{\prime}})}$ instead of $K_{\tau_{\theta}(\mathcal{P})}$. This approach enables the generation of cross-attention maps for all candidate classes within a single inference run.
@@ 3.4 Hierarchical Attention Refinement Method
In semantic segmentation tasks, attention maps play a crucial role in localizing and identifying different regions within an image. However, attention maps at a fixed resolution provide valuable localization information but may fail to capture the semantic diversity present across multiple spatial scales. As illustrated in Figure 5(a), cross-attention maps at various resolutions offer complementary insights: $\mathcal{C}^{8\times 8}$ provides coarse maps but accurate localization, while $\mathcal{C}^{16\times 16}$ captures better details of object position and shape. At higher resolutions (e.g., $\mathcal{C}^{32\times 32}$ and $\mathcal{C}^{64\times 64}$), attention maps may tend to degrade in quality and become increasingly complex to make sense of. These observations align with previous studies [44], which show that higher-resolution layers often focus on superficial features (such as color or style), while lower-resolution layers encode object semantics. To address this limitation, we are using a hierarchical cross-attention fusion mechanism that emphasizes semantically cross-attention maps (e.g., $\mathcal{C}^{16\times 16}$) while down-weighting noisy maps (e.g., $\mathcal{C}^{32\times 32}$ or $\mathcal{C}^{64\times 64}$). The fused cross-attention map $\widehat{\mathcal{C}}_{c}$ is computed as a weighted sum of the resolution-specific maps, each upsampled to a size of $64\times 64$:
@@
Each self-attention tensor $\mathcal{S}^{r}\in[0,1]^{r\times r\times r\times r}$ captures pairwise affinities among all spatial locations at resolution $r$. In a query-wise perspective, each spatial query position $(i,j)$ corresponds to a dedicated 2D attention map $\mathcal{S}^{r}_{[i,j,:,:]}\in[0,1]^{r\times r}$, which captures the semantic context relevant to that specific query. Figure 6 illustrates query-wise self-attention maps across different resolutions and spatial locations. The resolution of each query-wise attention map defines its receptive field in relation to the original image. To incorporate self-attention information into the segmentation process, we refine the fused cross-attention map $\widehat{\mathcal{C}_{c}}$ by applying affinity weighting derived from query-wise self-attention maps. Since $\widehat{\mathcal{C}_{c}}\in\mathbb{R}^{64\times 64}$ is defined at a fixed spatial resolution, we transform the original self-attention tensors $\mathcal{S}^{r}$ into $\mathcal{S}^{r\rightarrow 64}\in\mathbb{R}^{64\times 64\times 64\times 64}$ through upsampling and repetition. This transformation is illustrated in Figure 7, with an example from shape $2\times 2\times 2\times 2$ to $4\times 4\times 4\times 4$. While upsampling the last two dimensions is straightforward, repeating the first two dimensions relies on the empirical observation: 2D attention maps at neighboring positions often correspond to the same object group (e.g., $\mathcal{S}^{r}_{[i,j,:,:]}$ and $\mathcal{S}^{r}_{[i+1,j,:,:]}$, or $\mathcal{S}^{r}_{[i,j,:,:]}$ and $\mathcal{S}^{r}_{[i,j+1,:,:]}$) thus allowing spatial continuity to be preserved.
This transformation produces spatially aligned 4D self-attention tensors $\mathcal{S}^{r\rightarrow 64}$, which are compatible for multiplicative refinement of the class-specific cross-attention map $\widehat{\mathcal{C}}_{c}$. To incorporate self-attention affinities at multiple resolutions $r\in\mathcal{R}$, each $\mathcal{S}^{r\rightarrow 64}$ is matricized into an affinity matrix, and $\widehat{\mathcal{C}}_{c}$ is vectorized to match this dimension. These maps are then normalized and combined via a weighted sum over all resolutions to yield the final refined segmentation mask:
@@
By calculating $\widehat{\mathcal{A}}$ for the entire candidate class list $C$, we obtain the set of segmentation maps $\mathbf{A}=\{\widehat{\mathcal{A}}_{c}\mid c\in C\}$. Then, by applying the Test-Time Flipping (TTF) technique, we obtain two corresponding segmentation map lists: $\mathbf{A}^{\text{origin}}$ from the original image and $\mathbf{A}^{\text{flipped}}$ from the horizontally flipped image. We align the spatial information of the attention maps from the flipped image using $\mathcal{F}lip$ to match those from the original image. Then, we compute the average of each pair of corresponding attention maps from both lists to improve spatial consistency and prediction accuracy.
$\widehat{\mathcal{A}}_{c}^{*}=\frac{1}{2}\left[\widehat{\mathcal{A}}_{c}^{\text{origin}}+\mathcal{F}lip\Bigl{(}\widehat{\mathcal{A}}_{c}^{\text{flipped}}\Bigl{)}\right]$